# Subexponential lower bounds for $f$-ergodic Markov processes

### v1

@@ 3.1.1. Polynomial tails
[…]
@@
$\lim_{t\to\infty}\frac{\log\|\mathbb{P}_{x}(X_{t}\in\cdot)-\pi(\cdot))\|_{f_{k}}}{\log t}=\alpha_{k},\quad$ $\quad\lim_{t\to\infty}\frac{\log\mathbb{P}_{x}(\tau_{D}(\delta)>t)}{\log t}=m_{c}/(2-\ell),$
@@ 3.1.2. Stretched exponential tails
[…]
@@ 3.1.3. Exponential tails
[…]
@@ 3.2. Lévy-driven SDE
In this section we apply the results of Section 2 to a Lévy-driven Ornstein-Uhlenbeck process $X$ taking values in ${\mathcal{X}}={\mathbb{R}}$. We restrict ourselves to the additive noise case in order to keep the applications as concise as possible. We expect that, under appropriate asymptotic assumptions on the coefficients, the results of the present section can be generalised to the multi-dimensional setting with multiplicative Lévy noise.
Let $X$ be the solution of the following SDE,
(16) ${\mathrm{d}}X_{t}=-\mu X_{t}{\mathrm{d}}t+{\mathrm{d}}L_{t},$
where $\mu\in(0,\infty)$ and $L$ is a pure-jump Lévy process with Lévy measure $\nu$ (cf. [31, Ch. 1]). The drift and the Gaussian component of $L$ are assumed to be zero as they do not influence the asymptotic behaviour of $X$ when the Lévy measure $\nu$ has heavy tails. Adding them would require only a minor modification of the proof of Theorem 3.11 below.
@@
[…]
@@
Theorem 3.11 is applicable to general pure-jump Lévy drivers with two-sided jumps and arbitrary path variation. In particular, it covers the simpler model studied in [12], where it is assumed that $\nu((-\infty,1])=0$ (making $L$ a compound Poisson process with positive jumps and, consequently, ${\mathcal{X}}={\mathbb{R}}_{+}$ for $X_{0}=x\in{\mathbb{R}}_{+}$). Our lower bounds match the upper bounds from [12], i.e. for some constants $c_{\mathrm{TV}},C_{\mathrm{TV}}\in(0,\infty)$ we have
@@ 3.3. Stochastic damping Hamiltonian system
The upper bound on the rate of convergence in Theorem 3.14 is obtained by applying the drift condition in [9] (see (10) above) to an appropriate Lyapunov function (inspired by [34]). Note that the model in Theorem 3.14 has not been analysed in [9], where a hypoelliptic diffusion in (17) with stretched exponential tails is considered. The matching lower bound in Theorem 3.14 is obtained by comparing the tails of $X_{t}$ and $\pi$ via Lemma 5.1. This lower bound does not require a verification of the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$))., because the invariant measure $\pi$ of the process in (17) has a known density proportional to $(z,y)\mapsto\exp(-2c/\sigma^{2}(y^{2}/2+U(z)))$.
@@ 4. Return times to bounded sets for semimartingales
where the second inequality holds by the following facts: $\varphi$ is a non-decreasing function and the inequality $1/\kappa_{u}\leq 1/r$ is valid on the event $\{\rho_{r_{q}}<u<(\rho_{r_{q}}+t)\wedge\lambda_{r,\rho_{r_{q}}}\}$. The third inequality is a consequence of the fact $((\rho_{r_{q}}+t)\wedge\lambda_{r,\rho_{r_{q}}})-\rho_{r_{q}}=\tau_{1/r}\wedge t\leq t$, while the last eqaulity follows from the definition of $r_{q}$. By taking complements, we get
@@ 5.1. A lower bound on the $f$-variation rate of a Markov process
In this subsection we consider a strong Markov process $X=(X_{t})_{t\in{\mathbb{R}}_{+}}$ on a general metric space ${\mathcal{X}}$ with invariant measure $\pi$ on ${\mathcal{B}}({\mathcal{X}})$ (see Section 2.1 for definitions). The following lemma generalizes to $f$-variation the lower bound in [13, Thm 5.1] on the total variation between $\pi$ and the law of $X_{t}$. The key assumption in Lemma 5.1 is the lower bound on the decay of the tail of the integral of $f$ with respect to $\pi$. We stress that Lemma 5.1 does not require the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$))..
@@ 5.3. Proofs of the main results
By Assumption  (L($V$,$\varphi$,$\Psi$)). the process $X$ is positive Harris recurrent. The seminal result [25, Thm 1.2(b)] implies that a measurable $h:[1,\infty)\to[1,\infty)$ satisfies the implication:
@@
for some sufficiently large $r_{0}\in(0,\infty)$, where $G_{h}$ is the inverse of the increasing function $r\mapsto(1-q)h(r)/(2r\varphi(1/r))$. If the assumption in the implication in (29) holds, then the last integral in the previous display must also be infinite because the function $r\mapsto 1/\Psi(2G_{h}(r)/(1-q))$ is continuous and thus locally bounded. The criterion in [25, Thm 1.2(b)] stated above thus yields the conclusion of the implication in (29). ∎
@@ 6. Proofs of the examples in Section 3
This section is dedicated to proving the theorems presented in Section 3. All strong Markov processes $X$ considered in this section are ergodic with an invariant measure $\pi$ and Feller continuous (see [9, Sec. 5] for more details), and thus positive Harris recurrent by [25, Thm 1.1]. Moreover, all the models in this section are irreducible, either because of uniform ellipticity and the irreducibility of the driver (a Brownian motion or, more generally, a Lévy process) or by a direct argument in the hypoelliptic case. Thus, for any Lyapunov function $V$ for $X$, we will have $\mathbb{P}_{x}(V(X_{t})\geq r_{0})>0$ for all starting points $x$, times $t>0$ and levels $r_{0}$, thus satisfying the assumptions of Lemma A.1 below and implying non-confinement. Verifying the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$)). in our examples will thus reduce to finding functions $V,\varphi,\Psi$, which satisfy the conditions Assumption and Assumption in  (L($V$,$\varphi$,$\Psi$)).. Since we are working with Feller processes, we will obtain these conditions by applying the generator to the relevant functions, establishing the appropriate point-wise inequalities and using Theorem 2.7.
@@
Since the SDE in (11) possesses a unique strong solution, the process $X$ is strong Markov. Feller continuity follows from [16, Thm 3.4.1]. By Itô’s formula applied to $g(X)$, the extended generator (see Section 2.4 above for definition) of the diffusion $X$ takes the following form for any twice continuously differentiable $g\in C^{2}({\mathbb{R}}^{n})$:
@@
Let $X$ be the solution of SDE (16), driven by a pure-jump Lévy process $L$ with Lévy measure $\nu$. By Itô’s formula [26, Thm II.36] applied to $g(X)$, where $g\in C^{1}({\mathbb{R}})$ and $\int_{{\mathbb{R}}\setminus(-1,1)}|g(y)|\nu({\mathrm{d}}y)<\infty$, the extended generator (see Section 2.4 above for definition) of the process $X$ takes the form
(42) ${\mathcal{A}}g(x)=-\mu xg^{\prime}(x)+\int_{{\mathbb{R}}}(g(x+y)-g(x)-g^{\prime}(x)y{{\mathbbm{1}}\mkern-1.5mu}{\{y\in[-1,1]\}})\nu({\mathrm{d}}y)\qquad\text{for $x\in{\mathbb{R}}$.}$
@@
For any $g\in C^{2}({\mathbb{R}})$ and $C_{1}\coloneqq\int_{[-1,1]}y^{2}\nu({\mathrm{d}}y)\in[0,\infty)$, Lagrange’s theorem yields
(44) $\int_{[-1,1]}(g(x+y)-g(x)-g^{\prime}(x)y)\nu({\mathrm{d}}y)\leq C_{1}\sup_{u\in[x-1,x+1]}|g^{\prime\prime}(u)|\quad\text{for all $x\in{\mathbb{R}}$.}$
Part (a). Let $m\in[1,m_{c})$. For all large $x\in{\mathbb{R}}_{+}$ we have $g_{m}^{\prime}(x)=m(\log x)^{m-1}/x$, implying $g_{m}^{\prime}(x+u)\leq g_{m}^{\prime}(1+u)$ for all $u>0$. Thus, by Tonelli’s theorem, we have
$\int_{[1,\infty)}(g_{m}(x+y)-g_{m}(x))\nu({\mathrm{d}}y)$ $=\int_{[1,\infty)}\nu({\mathrm{d}}y)\int_{0}^{y}g_{m}^{\prime}(x+u){\mathrm{d}}u$
$=\int_{0}^{\infty}\nu([\max\{1,u\},\infty))g_{m}^{\prime}(x+u){\mathrm{d}}u$
(45) $\leq\int_{0}^{\infty}\nu([\max\{1,u\},\infty))g_{m}^{\prime}(1+u){\mathrm{d}}u<\infty,$
where the final inequality holds by 3.2 since $m<m_{c}$.
Since $g_{m}$ is non-decreasing and $\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|<\infty$ (by definition (43)), inequalities (44) and (45) imply there exists $x_{0}\in(0,\infty)$, such that the representation of ${\mathcal{A}}$ in (42) yields
(46) ${\mathcal{A}}g_{m}(x)$ $\leq-\mu m(\log x)^{m-1}+C_{1}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{[1,\infty)}(g_{m}(x+y)-g_{m}(x))\nu({\mathrm{d}}y)$
$\leq C_{1}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{0}^{\infty}\nu([\max\{1,u\},\infty))g_{m}^{\prime}(1+u){\mathrm{d}}u<\infty\qquad\text{for all $x\in[x_{0},\infty)$.}$
For $x\in(-\infty,x_{0})$, recall $\int_{[1,\infty)}g_{m}(x+u)\nu({\mathrm{d}}u)\leq\int_{[1,\infty)}g_{m}(x_{0}+u)\nu({\mathrm{d}}u)<\infty$ by (45). Hence the inequality in (44) and the global boundedness of $g_{m}^{\prime\prime}$ imply
${\mathcal{A}}g_{m}(x)\leq C_{1}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{[1,\infty)}g_{m}(x_{0}+u)\nu({\mathrm{d}}u)<\infty\quad\text{for all $x\in(-\infty,x_{0})$.}$
@@
$\int_{[-x/2,-1]}((1/g_{1})(x+y)-(1/g_{1})(x))\nu({\mathrm{d}}y)$ $=\int_{[-x/2,-1]}\left(\log(x)/\log(x+y)-1\right)\nu({\mathrm{d}}y)/\log(x)$
@@
${\mathcal{A}}(1/g_{1})(x)$ $\leq C_{1}\sup_{u\in[-1,1]}|(1/g_{1})^{\prime\prime}(x+u)|-\mu x(1/g_{1})^{\prime}(x)+\int_{(-\infty,-1]}\left(1/g_{1}(x+y)-1/g_{1}(x)\right)\nu({\mathrm{d}}y)$
$\leq C_{2}^{\prime}/g_{1}(x)^{2}+\nu((-\infty,-x/2])+\int_{[-x/2,-1]}((1/g_{1})(x+y)-(1/g_{1})(x))\nu({\mathrm{d}}y)$
$\leq C_{3}^{\prime}/g_{1}(x)^{2}\qquad\text{for all $x\geq x_{0}$and some constants $C_{2}^{\prime},C_{3}^{\prime}\in(0,\infty)$.}$
Jumps in $[1,\infty)$ can be disregarded since $1/g_{1}$ is non-increasing and the third inequality follows from 3.2 on the negative tail of $\nu$.
By (43), $1/g_{1}$ is non-increasing, $x(1/g_{1})^{\prime}(x)$ is bounded from below for $x\in{\mathbb{R}}$ and $1/g_{1}\leq 1$. Thus, since $\mu\in(0,\infty)$, we have
${\mathcal{A}}(1/g_{1})(x)\leq C_{1}\sup_{u\in{\mathbb{R}}}(1/g_{1})^{\prime\prime}(u)-\inf_{u\in{\mathbb{R}}}\mu u(1/g_{1})^{\prime}(u)+\nu((-\infty,-1])\eqqcolon b<\infty\quad\text{ for all $x\in{\mathbb{R}}$,}$
@@
(48) $X_{t}=x-\mu\int_{0}^{t}X_{s}{\mathrm{d}}s+L_{t}^{(-)}+L_{t}^{(M)}+L_{t}^{(+)},\quad\text{for all $t\in{\mathbb{R}}_{+}$,}$
@@
(49) $A_{t}\coloneqq\{L_{t}^{(-)}=0\}\cap\left\{\sup_{0\leq s\leq t}|L_{s}^{(M)}|\leq 1/8\right\}.$
@@
[…]
Evaluating the probability of the smallest event in the last display will complete the proof.
@@
Recall that $X$ follows (48) and the event $A_{t}$ in (49). Since the Lévy processes $L^{(-)}$ and $L^{(M)}$ are independent, started from zero and have right-continuous paths, for every $x\in{\mathbb{R}}$ we have
(50) $\mathbb{P}_{x}(A_{t})=\mathbb{P}(L_{t}^{(-)}=0)\>\mathbb{P}(\sup_{0\leq s\leq t}|L_{s}^{(M)}|\leq 1/8)>0\quad\text{ for all small $t\in(0,\infty)$.}$
Recall $V=g_{1}$, where $g_{1}$ is a $C^{2}({\mathbb{R}})$ function in (43). Choose $\ell_{0}\in(1,\infty)$ such that $V(y)=\log y$ for all $y\in[\ell_{0},\infty)$. Note that for all $\ell,r\in[\ell_{0},\infty)$, we have $S_{(\ell)}=\inf\{t>0:X_{t}<\exp(\ell)\}$ and $T^{(r)}=\inf\{t>0:X_{t}>\exp(r)\}$. Moreover, since $V^{-1}(r)=\exp(r)$, the set $\{V\geq r\}=[\exp(r),\infty)$ is an interval. We also assume that $\ell_{0}$ satisfies $\nu(\exp(\ell_{0}),\infty)<2\mu\exp(\ell_{0})$.
@@
Claim. For all $x\in\{V\geq\ell+1\}$ and $t\in(0,1/(2\mu\exp(\ell+2)))$ we have $\mathbb{P}_{x}(\{S_{(\ell)}\leq t\}\cap A_{t})=0$.
Proof of Claim. Pick $x\in\{V\geq\ell+1\}$. Since $\{\overline{\theta}_{k}\leq S_{(\ell)}<\underline{\theta}_{k}\}=\emptyset$ for all $k\in{\mathbb{N}}$, following holds

### v2

@@ 3.1.1. Polynomial tails
Assumption Ap. There exist $\alpha,\beta,\gamma\in(0,\infty)$ and $\ell\in[0,2)$, such that $2-\ell<2+(2\alpha-\gamma)/\beta=:m_{c}$ and the coefficients $b$ and $\Sigma=\sigma\sigma^{\intercal}$ of (11) satisfy (as $|x|\to\infty$)
@@
$\lim_{t\to\infty}\frac{\log\|\mathbb{P}_{x}(X_{t}\in\cdot)-\pi(\cdot))\|_{f_{k}}}{\log t}=\alpha_{k},$ $\quad\lim_{t\to\infty}\frac{\log\mathbb{P}_{x}(\tau_{D}(\delta)>t)}{\log t}=m_{c}/(2-\ell),$
@@ 3.1.2. Stretched exponential tails
Assumption Ase. There exist constants $p\in(0,1)$ and $\ell\in[0,2p)$, such that the coefficients $b$ and $\Sigma=\sigma\sigma^{\intercal}$ in (11) satisfy the following asymptotic assumptions: $\limsup_{|x|\to\infty}\operatorname{Tr}(\Sigma(x))/|x|^{\ell}<\infty$ and
@@ 3.1.3. Exponential tails
Assumption Ae. The coefficients $b$ and $\Sigma=\sigma\sigma^{\intercal}$ in (11) satisfy the following asymptotic assumptions: $\limsup_{|x|\to\infty}\operatorname{Tr}(\Sigma(x))<\infty$ and
@@ 3.2. Lévy-driven SDE
In this section we apply the results of Section 2 to a solution $X$ of a Lévy-driven SDE taking values in ${\mathcal{X}}={\mathbb{R}}$. More precisely, let $X$ follow the SDE
(16) ${\mathrm{d}}X_{t}=-\mu X_{t}{\mathrm{d}}t+\sigma(X_{t-}){\mathrm{d}}L_{t},$
where $\mu\in(0,\infty)$, $\sigma:{\mathbb{R}}\to{\mathbb{R}}_{+}$ is Lipschitz with $0<\inf_{x\in{\mathbb{R}}}\sigma(x)\leq\sup_{x\in{\mathbb{R}}}\sigma(x)<\infty$ and $L$ is a pure-jump Lévy process with Lévy measure $\nu$ (cf. [33, Ch. 1]). Such $X$ exists and is unique for any starting point $X_{0}=x\in{\mathbb{R}}$ [28, Thm V.6]. The drift and the Gaussian component of $L$ are zero as they do not influence the asymptotic behaviour of $X$ when $\nu$ has heavy tails and $\sigma$ is bounded. Adding them would require only a minor modification of the proof of Theorem 3.11 below.
@@
00 0 $<\liminf_{r\to\infty}(\log r)^{m_{c}}\nu([r,\infty))\leq\limsup_{r\to\infty}(\log r)^{m_{c}}\nu([r,\infty))<\infty\quad\text{and}$
@@
Theorem 3.11 is applicable to general pure-jump Lévy drivers with two-sided jumps and arbitrary path variation. In particular, it covers the simpler model studied in [12] with additive noise, where it is assumed that $\sigma\equiv 1$ and $\nu((-\infty,1])=0$ (making $L$ a compound Poisson process with positive jumps and, consequently, ${\mathcal{X}}={\mathbb{R}}_{+}$ for $X_{0}=x\in{\mathbb{R}}_{+}$). Our lower bounds match the upper bounds from [12], i.e. for some $c_{\mathrm{TV}},C_{\mathrm{TV}}\in(0,\infty)$ we have
@@ 3.3. Stochastic damping Hamiltonian system
The upper bound on the rate of convergence in Theorem 3.14 is obtained by applying the drift condition in [9] (see (10) above) to an appropriate Lyapunov function (inspired by [38]). Note that the model in Theorem 3.14 has not been analysed in [9], where a hypoelliptic diffusion in (17) with stretched exponential tails is considered. Following [15], the matching lower bound in Theorem 3.14 is obtained by comparing the tails of $X_{t}$ and $\pi$ via Lemma 5.1. This lower bound does not require a verification of the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$))., because the invariant measure $\pi$ of the process in (17) has a known density proportional to $(z,y)\mapsto\exp(-2c/\sigma^{2}(y^{2}/2+U(z)))$.
@@ 4. Return times to bounded sets for semimartingales
where the second inequality holds by the following facts: $\varphi$ is a non-decreasing function and the inequality $1/\kappa_{u}\leq 1/r$ is valid on the event $\{\rho_{r_{q}}<u<(\rho_{r_{q}}+t)\wedge\lambda_{r,\rho_{r_{q}}}\}$. The third inequality is a consequence of the fact $((\rho_{r_{q}}+t)\wedge\lambda_{r,\rho_{r_{q}}})-\rho_{r_{q}}=\tau_{1/r}\wedge t\leq t$, while the last equality follows from the definition of $r_{q}$. By taking complements, we get
@@ 5.1. A lower bound on the $f$-variation rate of a Markov process
In this subsection we consider a strong Markov process $X=(X_{t})_{t\in{\mathbb{R}}_{+}}$ on a general metric space ${\mathcal{X}}$ with invariant measure $\pi$ on ${\mathcal{B}}({\mathcal{X}})$ (see Section 2.1 for definitions). The following lemma generalizes to $f$-variation the lower bound in [15, Thm 3.6] on the total variation between $\pi$ and the law of $X_{t}$. The key assumption in Lemma 5.1 is the lower bound on the decay of the tail of the integral of $f$ with respect to $\pi$. We stress that Lemma 5.1 does not require the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$))..
@@ 5.3. Proofs of the main results
By Assumption  (L($V$,$\varphi$,$\Psi$)). the process $X$ is positive Harris recurrent. The seminal result [24, Thm 1.2(b)] implies that a measurable $h:[1,\infty)\to[1,\infty)$ satisfies the implication:
@@
for some sufficiently large $r_{0}\in(0,\infty)$, where $G_{h}$ is the inverse of the increasing function $r\mapsto(1-q)h(r)/(2r\varphi(1/r))$. If the assumption in the implication in (29) holds, then the last integral in the previous display must also be infinite because the function $r\mapsto 1/\Psi(2G_{h}(r)/(1-q))$ is continuous and thus locally bounded. The criterion in [24, Thm 1.2(b)] stated above thus yields the conclusion of the implication in (29). ∎
@@ 6. Proofs of the examples in Section 3
This section is dedicated to proving the theorems presented in Section 3. All strong Markov processes $X$ considered in this section are ergodic with an invariant measure $\pi$ and Feller continuous (see [9, Sec. 5] for more details), and thus positive Harris recurrent by [24, Thm 1.1]. Moreover, all the models in this section are irreducible, either because of uniform ellipticity and the irreducibility of the driver (a Brownian motion or, more generally, a Lévy process) or by a direct argument in the hypoelliptic case. Thus, for any Lyapunov function $V$ for $X$, we will have $\mathbb{P}_{x}(V(X_{t})\geq r_{0})>0$ for all starting points $x$, times $t>0$ and levels $r_{0}$, thus satisfying the assumptions of Lemma A.1 below and implying non-confinement. Verifying the $\mathbf{L}$-drift condition  (L($V$,$\varphi$,$\Psi$)). in our examples will thus reduce to finding functions $V,\varphi,\Psi$, which satisfy the conditions Assumption and Assumption in  (L($V$,$\varphi$,$\Psi$)).. Since we are working with Feller processes, we will obtain these conditions by applying the generator to the relevant functions, establishing the appropriate point-wise inequalities and using Theorem 2.7.
@@
Since the SDE in (11) possesses a unique strong solution, the process $X$ is strong Markov. Feller continuity follows from [17, Thm 3.4.1]. By Itô’s formula applied to $g(X)$, the extended generator (see Section 2.4 above for definition) of the diffusion $X$ takes the following form for any twice continuously differentiable $g\in C^{2}({\mathbb{R}}^{n})$:
@@
Let $X$ be the solution of SDE (16) with a bounded Lipschitz dispersion coefficient $\sigma$, driven by a pure-jump Lévy process $L$ with Lévy measure $\nu$. By Itô’s formula [28, Thm II.36] applied to $g(X)$, where $g\in C^{1}({\mathbb{R}})$ and $\int_{{\mathbb{R}}\setminus(-1,1)}|g(y)|\nu({\mathrm{d}}y)<\infty$, the extended generator (see Section 2.4 above for definition) of the process $X$ takes the form
(42) ${\mathcal{A}}g(x)=-\mu xg^{\prime}(x)+\int_{{\mathbb{R}}}(g(x+\sigma(x)y)-g(x)-g^{\prime}(x)\sigma(x)y{{\mathbbm{1}}\mkern-1.5mu}{\{y\in[-1,1]\}})\nu({\mathrm{d}}y),\qquad\text{$x\in{\mathbb{R}}$.}$
@@
For any $g\in C^{2}({\mathbb{R}})$ and $C_{1}^{\prime}\coloneqq\sup_{x\in{\mathbb{R}}}\sigma(x)^{2}\int_{[-1,1]}y^{2}\nu({\mathrm{d}}y)\in[0,\infty)$, Lagrange’s theorem yields
(44) $\int_{[-1,1]}(g(x+\sigma(x)y)-g(x)-g^{\prime}(x)\sigma(x)y)\nu({\mathrm{d}}y)\leq C_{1}^{\prime}\sup_{u\in[x-\sigma(x),x+\sigma(x)]}|g^{\prime\prime}(u)|\quad\text{for all $x\in{\mathbb{R}}$.}$
Part (a). Let $m\in[1,m_{c})$. For all large $x\in{\mathbb{R}}_{+}$ we have $g_{m}^{\prime}(x)=m(\log x)^{m-1}/x$, implying $g_{m}^{\prime}(x+u)\leq g_{m}^{\prime}(1+u)$ for all $u>0$. Thus, by Tonelli’s theorem, for all large $x\in{\mathbb{R}}_{+}$ we have
$\int_{[1,\infty)}(g_{m}(x+\sigma(x)y)-g_{m}(x))\nu({\mathrm{d}}y)$ $=\int_{[1,\infty)}\nu({\mathrm{d}}y)\int_{0}^{\sigma(x)y}g_{m}^{\prime}(x+u){\mathrm{d}}u$
$=\int_{0}^{\infty}\nu([\max\{1,u/\sigma(x)\},\infty))g_{m}^{\prime}(x+u){\mathrm{d}}u$
(45) $\leq\int_{0}^{\infty}\nu([\max\{1,u/\sup_{y\in{\mathbb{R}}}\sigma(y)\},\infty))g_{m}^{\prime}(1+u){\mathrm{d}}u<\infty,$
where the final inequality holds by 3.2 since $m<m_{c}$ and $\sigma$ is bounded on ${\mathbb{R}}$.
Since $g_{m}$ is non-decreasing and $\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|<\infty$ (by definition (43)), inequalities (44) and (45) imply there exists $x_{0}\in(0,\infty)$, such that, for all $x\in[x_{0},\infty)$, the representation of ${\mathcal{A}}$ in (42) yields
(46) ${\mathcal{A}}g_{m}(x)$ $\leq-\mu m(\log x)^{m-1}+C_{1}^{\prime}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{[1,\infty)}(g_{m}(x+\sigma(x)y)-g_{m}(x))\nu({\mathrm{d}}y)$
$\leq C_{1}^{\prime}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{0}^{\infty}\nu([\max\{1,u/\sup_{y\in{\mathbb{R}}}\sigma(y)\},\infty))g_{m}^{\prime}(1+u){\mathrm{d}}u<\infty.$
Let $x\in(-\infty,x_{0})$. Then $\int_{[1,\infty)}g_{m}(x+\sigma(x)y)\nu({\mathrm{d}}y)\leq\int_{[1,\infty)}g_{m}(x_{0}+\sup_{u\in{\mathbb{R}}}\sigma(u)y)\nu({\mathrm{d}}y)<\infty$ since $g_{m}$ is non-decreasing and integrable with respect to $\nu$ by 3.2. Hence the inequality in (44) and the global boundedness of $g_{m}^{\prime\prime}$ imply
${\mathcal{A}}g_{m}(x)\leq C_{1}^{\prime}\sup_{u\in{\mathbb{R}}}|g_{m}^{\prime\prime}(u)|+\int_{[1,\infty)}g_{m}(x_{0}+\sup_{y\in{\mathbb{R}}}\sigma(y)u)\nu({\mathrm{d}}u)<\infty\quad\text{for all $x\in(-\infty,x_{0})$.}$
@@
$\int_{[-\frac{x}{2\sigma(x)},-1]}((1/g_{1})(x+\sigma(x)y)-(1/g_{1})(x))\nu({\mathrm{d}}y)$ $=\int_{[-\frac{x}{2\sigma(x)},-1]}\frac{\log x-\log(x+\sigma(x)y)}{\log(x+\sigma(x)y)\log x}\nu({\mathrm{d}}y)$
@@
${\mathcal{A}}(1/g_{1})(x)$ $\leq C_{1}^{\prime}\sup_{u\in[x-\sigma(x),x+\sigma(x)]}|(1/g_{1})^{\prime\prime}(u)|-\mu x(1/g_{1})^{\prime}(x)+\int_{(-\infty,-1]}\frac{g_{1}(x)-g_{1}(x+\sigma(x)y)}{g_{1}(x)g_{1}(x+\sigma(x)y)}\nu({\mathrm{d}}y)$
$\leq C_{2}^{\prime}/g_{1}(x)^{2}+\nu((-\infty,-\frac{x}{2\sigma(x)}))+\int_{[-\frac{x}{2\sigma(x)},-1]}((1/g_{1})(x+\sigma(x)y)-(1/g_{1})(x))\nu({\mathrm{d}}y)$
$\leq C_{2}/g_{1}(x)^{2}\qquad\text{for all $x\geq x_{0}$ and someconstants $C_{2}^{\prime},C_{2}\in(0,\infty)$.}$
Jumps in $[1,\infty)$ can be disregarded since $1/g_{1}$ is non-increasing and the third inequality follows from 3.2 on the negative tail of $\nu$, bound (47) and the fact that $\sigma$ is bounded.
By (43), $1/g_{1}$ is non-increasing, $x(1/g_{1})^{\prime}(x)$ is bounded from below for $x\in{\mathbb{R}}$ and $1/g_{1}\leq 1$. Thus, since $\mu\in(0,\infty)$, by (44) we have
${\mathcal{A}}(1/g_{1})(x)\leq C_{1}^{\prime}\sup_{u\in{\mathbb{R}}}(1/g_{1})^{\prime\prime}(u)-\inf_{u\in{\mathbb{R}}}\mu u(1/g_{1})^{\prime}(u)+\nu((-\infty,-1])\eqqcolon b<\infty\quad\text{ for all $x\in{\mathbb{R}}$,}$
@@
(48) $X_{t}=x-\mu\int_{0}^{t}X_{s}{\mathrm{d}}s+\int_{0}^{t}\sigma(X_{s-}){\mathrm{d}}L_{s}^{(-)}+\int_{0}^{t}\sigma(X_{s-}){\mathrm{d}}L_{s}^{(M)}+\int_{0}^{t}\sigma(X_{s-}){\mathrm{d}}L_{s}^{(+)},\quad\text{for all $t\in{\mathbb{R}}_{+}$,}$
@@
(49) $A_{t}\coloneqq\{L_{t}^{(-)}=0\}\cap\left\{\sup_{0\leq s\leq t}\left|\int_{0}^{s}\sigma(X_{u-}){\mathrm{d}}L_{u}^{(M)}\right|\leq 1/8\right\}.$
@@
$\{T^{(r)}<S_{(\ell)}\}\supset\{T^{(r)}<S_{(\ell)}\}\cap A_{t}\supset\{T^{(r)}<t\}\cap A_{t}\supset\left\{\sigma_{-}\sup_{0\leq s\leq t}(L_{s}-L_{s-})\geq V^{-1}(r)\right\}\cap A_{t}$
$\mathbb{P}_{x}$-a.s., where $\sigma_{-}\coloneqq\inf_{u\in{\mathbb{R}}}\sigma(u)>0$. Evaluating the probability of the smallest event in the last display will complete the proof.
@@
Recall that $X$ follows (48) and $V=g_{1}$, where $g_{1}$ is a $C^{2}({\mathbb{R}})$ function in (43). Choose $\ell_{0}\in(1,\infty)$ such that $V^{-1}(y)=\exp(y)$ for all $y\in[\ell_{0},\infty)$. Note that for all $\ell,r\in[\ell_{0},\infty)$, we have $S_{(\ell)}=\inf\{t>0:X_{t}<\exp(\ell)\}$ and $T^{(r)}=\inf\{t>0:X_{t}>\exp(r)\}$. By possibly increasing $\ell_{0}$, we may (and do) assume that $\nu(\exp(\ell_{0})/\sigma_{-},\infty)<2\mu\exp(\ell_{0})$ holds.
@@
Claim 1. For all $x\in\{V\geq\ell+1\}$ and $t\in(0,1/(2\mu\exp(\ell+2)))$ we have $\mathbb{P}_{x}(\{S_{(\ell)}\leq t\}\cap A_{t})=0$.
Proof of Claim 1. Pick $x\in\{V\geq\ell+1\}$. Since $\{\overline{\theta}_{k}\leq S_{(\ell)}<\underline{\theta}_{k}\}=\emptyset$ for all $k\in{\mathbb{N}}$, the following holds